OpenStudy (anonymous):

**PLEASE HELP, I'M DESPERATE *** Brayden is investing $1,617 in three different stocks: Zycodec, Unix, and Xenon. He will invest 40% of the total between Zycodec and Xenon in the Unix stock. Twice as much as the investment in Zycodec will be given to Xenon. How much will Brayden invest in the Unix stock? A. $385 B. $462 C. $770 D. $924

OpenStudy (zzr0ck3r):

I don't understand this line "He will invest 40% of the total between Zycodec and Xenon in the Unix stock"

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

OMGGGGGG IHATE MATHHHHHHHHHH

OpenStudy (jdoe0001):

lol

OpenStudy (zzr0ck3r):

are you sure those are the exact words?>

OpenStudy (anonymous):

Yes, why wouldn't they be? I copied and pasted

OpenStudy (jdoe0001):

well, I get this Z + X = SOME AMOUNT U = 40% SOME AMOUNT => U = (Z+X)0.4 now twice as much of Z goes to X so X = 2Z

OpenStudy (zzr0ck3r):

ahh ok I see what its saying

OpenStudy (jdoe0001):

so Z + X + U = 1617 but X = \(\color{blue}{2Z}\) and U = (Z + X) => ( Z + \(\color{blue}{2Z}\))

OpenStudy (jdoe0001):

hmm, actually and U = (Z + X) => ( Z + \(\color{blue}{2Z}\))(0.4)

OpenStudy (jdoe0001):

which is 40% of the total of Zycoden and Xenon stocks

OpenStudy (jdoe0001):

so Z + X + U = 1617 or Z + 2Z + (Z + 2Z)(0.4) = 1617 then just solve for "Z"

OpenStudy (anonymous):

omfg

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

z + 2z + (z+2z)(0.4)

OpenStudy (anonymous):

Lower case makes it better but I still don't understand

OpenStudy (jdoe0001):

once you find Z, zycodec amount to get Unix stock amount, well, U = (Z + 2Z)0.4

OpenStudy (anonymous):

Idk how to find z

OpenStudy (jdoe0001):

is just plain simplifying

OpenStudy (jdoe0001):

the Xenon stock is based on the Zycodec, 2 to 1, so Xenon's is twice as much or 2Z the total amount invested in Zycodec and Xenon is , Zycodec's plus Xenon or Z + 2Z

OpenStudy (jdoe0001):

and the Unix stock is based on their total, 40% of it

OpenStudy (anonymous):

B. $462 ????

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

Zycodec's amount is $385 thus 0.4 or 2/5 2/5 ( 385 + 770) => 462